# Critical behaviour of the fully packed loop-$O(n)$ model on planar triangulations

### v1

@@ Abstract
We study the fully packed loop $O(n)$ model on planar triangulations. This model is also bijectively equivalent to the Fortuin–Kasteleyn model of planar maps with parameter $q\in(0,4)$ at its self-dual point. These have been traditionally studied using either techniques from analytic combinatorics (based in particular on the gasket decomposition of Borot, Bouttier and Guitter [BBG12a]) or probabilistic arguments (based on Sheffield’s “hamburger-cheeseburger” bijection [sheffield2016quantum]). In this paper we establish a dictionary relating quantities of interest in both approaches. This has several consequences. First, we derive an exact expression for the partition function of the fully packed loop-$O(n)$ model on triangulations, as a function of the outer boundary length. This confirms predictions by Gaudin and Kostov [gaudin1989n]. In particular, this model exhibits critical behaviour, in the sense that the partition function exhibits a power-law decay characteristic of the critical regime at this self-dual point. This can be thought of as the planar map analogue of Nienhuis’ predictions [Nienhuis82, Nienhuis84] for the critical point of the loop-$O(n)$ model on the hexagonal lattice. Finally, we derive precise asymptotics for geometric features of the FK model of planar maps when $0<q<4$, such as the exact tail behaviour of the perimeters of clusters and loops. This sharpens previous results of [berestycki2017critical] and [gwynne2019scaling]. A key step is to use the above dictionary and the probabilistic results to justify rigorously an ansatz commonly assumed in the analytic combinatorics literature.
@@ 1 Introduction
The second model we will discuss is that of the fully packed loop $O(n)$ model on planar triangulations. In this model we sample a triangulation $T$, together with a fully packed configuration $L$ of loops, with probability
@@
In many models of planar statistical mechanics, it is widely expected that the self-dual point coincides with its critical point (see for instance [grimmett2018probability] for some background discussion). In particular, such a statement is the content of a celebrated result of Beffara and Duminil-Copin [beffara2012self] for the random cluster model (equivalently the Fortuin–Kasteleyn percolation model) on the square lattice and for $q\geq 1$. For the related loop $O(n)$ model on the hexagonal lattice, the equality between critical and self-dual point led Nienhuis to predict ([Nienhuis82, Nienhuis84]) that the critical point occurs at $x=x_{c}=1/\sqrt{2+\sqrt{2-n}}$ for $n\in(0,2]$, and in fact $n\in[-2,2]$; this remains a famous open problem in this field (see [DCensaios] for a thorough survey). This should also apply to the two self-dual models planar maps discussed above, (1.2) and (1.3). This suggests that this model is also critical, although in a sense that needs to be carefully specified.
Furthermore, one of the principal conjectures in the field – closely related to the DDK ansatz mentioned above – states that, when suitably renormalised and conformally embedded into the Riemann sphere, self-dual FK(q)(q)-weighted planar maps converge to a $\gamma$-LQG surface, where
@@
On the other hand, a classical approach to the loop $O(n)$ model is to make use of the spatial Markov property of the model. This results in the so-called gasket decomposition which was in particular used by Borot, Bouttier and Guitter [BBG12, BBG12a, BBG12b] to establish the phase diagram of the loop-$O(n)$ model. Building on powerful tools of analytic combinatorics they derived and solved (assuming a certain natural ansatz) an equation for the resolvent of the partition function of the model, which yields fine asymptotics. Such an approach has been made fully rigorous (including the proof of the above ansatz) by works of Budd and Chen [budd2019peeling], in the case of the so-called rigid model on quadrangulations (where loops are constrained to enter and leave a given quadrangle through opposite edges). From this, Chen, Curien and Maillard [ChenCurienMaillard] deduced some scaling limit results for the perimeter cascade of $O(n)$ loops, and Aïdékon, Da Silva and Hu [aidekon2024scaling] proved the scaling limit for the volume of such rigid loop-$O(n)$ quadrangulations. A similar approach was also used by Borot, Duplantier and Guitter [BBD16] to determine the nesting statistics in the bending energy variant of the $O(n)$ model, where loops can bend inside a quadrangle at a given energetic cost.
@@
We stress that this dictionary has been used in [da2025scaling], which concerns the critical case $q=4$ (equivalently, $n=2$). In that case, the authors prove the analogue of 1.1 by directing solving the gasket decomposition equation (without any extra input) and then deduce the scaling limit of FK(4)(4)-weighted planar maps in the peanosphere sense, resolving the remaining open regime in Sheffield’s work [sheffield2016quantum].
@@ Acknowledgments.
We thank Sasha Glazman, Xingjian Hu, Ellen Powell, Xin Sun, Joonas Turunen and Mo Dick Wong for insightful discussions. N.B. acknowledges the support from the Austrian Science Fund (FWF) grants 10.55776/F1002 on “Discrete random structures: enumeration and scaling limits” and 10.55776/PAT1878824 on “Random Conformal Fields”. W.D.S. is supported by the Austrian Science Fund (FWF) grant on “Emergent branching structures in random geometry” (DOI: 10.55776/ESP534).
@@ 3.3 Dictionary and proof of ansatz
We now have the tools we need in order to express our correspondence between loop $O(n)$ statistics and FK planar maps. The correspondence takes the form of an exact identity between the partition function $F_{\ell}$ of the loop $O(n)$ model with boundary size $\ell$, and hitting times of the reduced (burger) walk. This will essentially follow from summing 3.4 over all maps in $\mathbb{T}_{\ell}$ and the connection that we made in 3.2 between the reduced walk and the typical filled-in cluster.
@@ Appendix A Computation of inverse Fourier transform
$\int_{0}^{\infty}\frac{\cosh(2bt)}{(\cosh t)^{2a}}\mathrm{d}t=4^{a-1}B(a+b,a-b),$ (A.2)

### v2

@@ Abstract
We study the fully packed loop-$O(n)$ model on planar triangulations. This model is also bijectively equivalent to the Fortuin–Kasteleyn model of planar maps with parameter $q\in(0,4)$ at its self-dual point. These have been traditionally studied using either techniques from analytic combinatorics (based in particular on the gasket decomposition of Borot, Bouttier and Guitter [BBG12a]) or probabilistic arguments (based on Sheffield’s “hamburger-cheeseburger” bijection [sheffield2016quantum]). In this paper we establish a dictionary relating quantities of interest in both approaches. This has several consequences. First, we derive an exact expression for the partition function of the fully packed loop-$O(n)$ model on triangulations, as a function of the outer boundary length. This confirms predictions by Gaudin and Kostov [gaudin1989n]. In particular, this model exhibits critical behaviour, in the sense that the partition function exhibits a power-law decay characteristic of the critical regime at this self-dual point. Finally, we derive precise asymptotics for geometric features of the FK model of planar maps when $0<q<4$, such as the exact tail behaviour of the perimeters of clusters and loops. This sharpens previous results of [berestycki2017critical] and [gwynne2019scaling]. A key step is to use the above dictionary and the probabilistic results to justify rigorously an ansatz commonly assumed in the analytic combinatorics literature.
@@ 1 Introduction
The second model we will discuss is that of the fully packed loop-$O(n)$ model on planar triangulations. In this model we sample a triangulation $T$, together with a fully packed configuration $L$ of loops, with probability
@@
Both models are believed to be critical in some sense, as we now explain.
On the one hand, from the perspective of the self-dual FK model (1.2), criticality (if it indeed holds) would agree with the fact that, in common with many planar models of statistical mechanics, the self-dual point coincides with its transition point (see for instance [grimmett2018probability] for some background discussion). In particular, such a statement is the content of a celebrated result of Beffara and Duminil-Copin [beffara2012self] for the random cluster model (equivalently the Fortuin–Kasteleyn percolation model) on the square lattice and for $q\geq 1$.
[…]
$z^{|\boldsymbol{\ell}|}n^{\#\boldsymbol{\ell}},$ (1.5)
[…]
On the hexagonal lattice, the loop-$O(n)$ model (1.5) is predicted to have a rich phase diagram (see e.g. [peled2019lectures, Section 3]). Notably, it was predicted by Nienhuis [Nienhuis82, Nienhuis84] that there is a critical line, given by
$z=z_{c}(n)=1/\sqrt{2+\sqrt{2-n}}$
for $n\in(0,2]$ (and in fact $n\in[-2,2]$ in the physics literature) separating the sub-critical phase ($z<z_{c}$) where the loops exhibit exponential decay and the critical phase ($z\geq z_{c}$) where the decay should only be polynomial. In fact, for fixed $n$, there should be two “critical” regimes called dilute and dense respectively, according as $z=z_{c}(n)$ or $z>z_{c}(n)$, where the model is expected to have different (but both conformally-invariant) scaling limits. See [kager2004guide, Section 5.6] for a statement of this conjecture. Although this remains a famous open problem in the field, we also stress that spectacular progress has been made in various regions of the phase diagram – we refer to the recent paper [glazman2025planar] for the state of the art and an important breakthrough in this direction. It is not altogether clear whether the fully-packed case on the hexagonal lattice should correspond to a third phase with a distinct conformally invariant scaling limit or should be in the same universality class as the dense phase (see [blote1994fully] for some discussion). Nonrigorous renormalisation arguments have suggested that the fully-packed case corresponds to an unstable fixed (critical) point, although on the square lattice, according to some predictions [blote2012completely] the fully packed model appears to behave as a dense phase. In the setting of planar maps, there seems to be a consensus that the fully-packed case belongs to the same universality class as the dense case and our results below support this. It would be interesting to study the fully packed model further on the hexagonal and square lattices, since typically, one expects the same behaviour on a fixed lattice and a random one when there is universality and/or conformal invariance.
Turning back to the predictions concerning planar maps, in both cases, one of the principal conjectures in the field – closely related to the DDK ansatz mentioned above – states that, when suitably renormalised and conformally embedded into the Riemann sphere, self-dual FK(q)(q)-weighted planar maps and fully packed loop-$O(n)$ decorated planar maps converge to a $\gamma$-LQG surface, where
@@
On the other hand, a classical approach to the loop-$O(n)$ model is to make use of the spatial Markov property of the model. This results in the so-called gasket decomposition which was in particular used by Borot, Bouttier and Guitter [BBG12, BBG12a, BBG12b] to establish the phase diagram of the loop-$O(n)$ model. Building on powerful tools of analytic combinatorics they derived and solved (assuming a certain natural ansatz) an equation for the resolvent of the partition function of the model, which yields fine asymptotics. Such an approach has been made fully rigorous (including the proof of the above ansatz) by works of Budd and Chen [budd2019peeling], in the case of the so-called rigid model on quadrangulations (where loops are constrained to enter and leave a given quadrangle through opposite edges). From this, Chen, Curien and Maillard [ChenCurienMaillard] deduced some scaling limit results for the perimeter cascade of $O(n)$ loops, and Aïdékon, Da Silva and Hu [aidekon2024scaling] proved the scaling limit for the volume of such rigid loop-$O(n)$ quadrangulations. A similar approach was also used by Borot, Duplantier and Guitter [BBD16] to determine the nesting statistics in the bending energy variant of the $O(n)$ model, where loops can bend inside a quadrangle at a given energetic cost.
@@
We stress that this dictionary has been used in the concurrent work [da2025scaling], which concerns the critical case $q=4$ (equivalently, $n=2$). In that case, the authors prove the analogue of 1.1 by directing solving the gasket decomposition equation (without any extra input) and then deduce the scaling limit of FK(4)(4)-weighted planar maps in the peanosphere sense, resolving the remaining open regime in Sheffield’s work [sheffield2016quantum].
@@ Acknowledgments.
We thank Xingjian Hu, Ellen Powell, Xin Sun, Joonas Turunen and Mo Dick Wong for insightful discussions. We are especially grateful to Sasha Glazman for discussions concerning the literature on the loop $O(n)$ model on the hexagonal lattice, and helping us clarify some confusions regarding this in a preliminary version of the paper. N.B. acknowledges the support from the Austrian Science Fund (FWF) grants 10.55776/F1002 on “Discrete random structures: enumeration and scaling limits” and 10.55776/PAT1878824 on “Random Conformal Fields”. W.D.S. is supported by the Austrian Science Fund (FWF) grant on “Emergent branching structures in random geometry” (DOI: 10.55776/ESP534).
@@ 3.3 Dictionary and proof of ansatz
We now have the tools we need in order to express our correspondence between loop-$O(n)$ statistics and FK planar maps. The correspondence takes the form of an exact identity between the partition function $F_{\ell}$ of the loop-$O(n)$ model with boundary size $\ell$, and hitting times of the reduced (burger) walk. This will essentially follow from summing 3.4 over all maps in $\mathbb{T}_{\ell}$ and the connection that we made in 3.2 between the reduced walk and the typical filled-in cluster.
@@ Appendix A Computation of inverse Fourier transform
$\int_{0}^{\infty}\frac{\cosh(2bt))}{(\cosh t)^{2a}}\mathrm{d}t=4^{a-1}B(a+b,a-b),$ (A.2)